OpenStudy (dj3strella):

Question 4.4. Transform this equation to slope-intercept form: 6x + 4y = 8. (Points : 1)

OpenStudy (anonymous):

6x + 4y = 8 6x = 8 - 4y 1.5x = 2 - y Answer: y = -1.5x - 2

OpenStudy (anonymous):

BTW, the multiple equations above the answer are the steps I took to solve it.